# HyperMoE: Towards Better Mixture of Experts via Transferring Among Experts

### v1

@@ Abstract
The Mixture of Experts (MoE) for language models has been proven effective in augmenting the capacity of models by dynamically routing each input token to a specific subset of experts for processing. Despite the success, most existing methods face a challenge for balance between sparsity and the availability of expert knowledge: enhancing performance through increased use of expert knowledge often results in diminishing sparsity during expert selection. To mitigate this contradiction, we propose HyperMoE, a novel MoE framework built upon Hypernetworks. This framework integrates the computational processes of MoE with the concept of knowledge transferring in multi-task learning. Specific modules generated based on the information of unselected experts serve as supplementary information, which allows the knowledge of experts not selected to be used while maintaining selection sparsity. Our comprehensive empirical evaluations across multiple datasets and backbones establish that HyperMoE significantly outperforms existing MoE methods under identical conditions concerning the number of experts. We provide some weights for HyperMoE at https://github.com/Bumble666/HyperMoE_early_version
@@ 1 Introduction
The accelerated advancement of large language models has culminated in their widespread application across various domains, including healthcare, education, and social interactions (Brown et al.,, 2020; Achiam et al.,, 2023; Touvron et al.,, 2023). The remarkable capabilities of these models are attributed to the enhancements in their scale. Nevertheless, the scaling of dense models is often hampered by significant computational demands, posing a challenge to developing the Natural Language Processing (NLP) community. In response, sparse activation models have emerged as a solution (Artetxe et al.,, 2022; Du et al.,, 2022), activating only a subset of parameters for different inputs, thus mitigating computational costs. One of the most representative methods is the Mixture of Experts (MoE, Shazeer et al., (2017)), which routers different inputs to specific groups of experts, thereby enlarging the model’s capacity without increasing computational burdens.
@@
The key to effectively reducing computational costs lies in the sparsity of expert selection, with the number of experts selected for each token being kept at a lower level. In practical applications or experiments, existing works (Roller et al., 2021a, ; Fedus et al.,, 2022; Rajbhandari et al.,, 2022; Xue et al.,, 2023) usually select only one or two experts per input. However, increasing the number of selected experts per token can enhance the availability of expert knowledge and improve the performance of downstream tasks Yang et al., (2019); Shazeer et al., (2017); He et al., (2023). This scenario positions the MoE model in a predicament akin to a zero-sum game: a choice between increasing the number of available experts to improve performance or preserving a lower level of available experts to ensure sparsity, as depicted in Figure 1.
To mitigate this contradiction, one solution would be to use the knowledge of other experts to assist the sparsely selected experts. This is similar to multi-task learning, which transfers knowledge among related tasks. Some works (Karimi Mahabadi et al.,, 2021; Ivison and Peters,, 2022; Zhao et al.,, 2023) suggest using hypernetworks (Ha et al.,, 2017) to generate task-specific knowledge to enhance positive transfer between tasks. Inspired by this, we aim to increase the availability of expert knowledge by transferring the knowledge of unselected experts while sparsely selecting experts.
@@
We further improve upon this by introducing the concept of cross-layer Hypernetworks: A hypernetwork is shared among all transformer layers, which enables information flow among MoEs in different layers. This brings additional efficiency in terms of parameters and computational costs: Despite the additional computation, our method only experienced a decrease [^1] The degree of decline in speed is related to the scale of the Hypernetworks and the bottleneck size in the generated HyperExpert (similar to $r$ in LoRA (Hu et al.,, 2022)). For various tasks, these hyperparameters can be dynamically adjusted to control the delay. of approximately 15% in training speed and 10% in inference speed compared to the standard MoE.
We evaluate HyperMoE on 20 representative NLP datasets across diverse tasks: sequence classification, extractive question answering, summarization, and text generation. Extensive experimental results show that HyperMoE outperforms baselines, including Switch Transformer (Fedus et al.,, 2022) with MoE architecture. This demonstrates the effectiveness of our method in transferring knowledge to experts, which increases the utilization of expert knowledge while keeping the number of experts selected at a low level.
@@
  • We propose a novel HyperMoE architecture with HyperExpert for the MoE framework, which resolves the inherent tension between maintaining sparse expert selection and ensuring sufficient expert availability within MoE.
  • HyperMoE outperforms strong baselines based on Switch Transformer across a diverse set of NLP tasks, confirming our approach’s effectiveness.
@@ 2.1 Mixture of Expert
The gate model $G(\cdot)$ is a Noisy Top-K Network Shazeer et al., (2017) with parameters $W_{g}$ and $W_{noise}$. This gating method introduces adjustable noise and then retains the top-k values as the final output:
@@ 2.2 HyperNetworks
Hypernetwork (Ha et al.,, 2017) can generate parameters to be used by target networks or modules. Specifically, a hypernetwork with independent parameters $\phi$ denoted as $h_{\phi}$, leverages an context information $z$ to generate the target parameters $\theta$ for the primary network $f_{\theta}$ and the primary network with an input $x$ is redefined as:
@@
This method of flexibly adjusting the parameters of the target network to adapt to different input scenarios is widely used in multi-task learning (Karimi Mahabadi et al.,, 2021; Üstün et al.,, 2022) and few-shot learning (Ponti et al.,, 2021). While generating condition-specific parameters, these parameters also benefit from shared knowledge (Pfeiffer et al.,, 2023).
@@ 3.1 Conditional Expert
In our approach, the matrices $W_{1}$ and $W_{2}$ are generated by a hypernetwork as described in Section 3.2. In addition, we adopt a bottleneck structure for the conditional expert to improve parameter efficiency inspired by the Adapter (Houlsby et al.,, 2019). Specifically, the bottleneck dimension $b$ satisfies $b\ll h$ in our method.
@@ 3.2 HyperExpert
These works (Karimi Mahabadi et al.,, 2021; He et al.,, 2022; Phang et al.,, 2023; Ivison et al.,, 2023) indicate that hypernetworks can learn the parameter information of the main neural network under different input scenarios and efficiently adjust the parameters of the target network to adapt to this information.
@@
The weight matrices of hypernetworks ${W^{D/U}}$ are used to generate the down-projection matrix ${D}_{i}^{\tau}\in\mathbb{R}^{h\times b}$ and the up-projection matrix ${U}_{i}^{\tau}\in\mathbb{R}^{b\times h}$ in the HyperExpert $\hat{E}_{i}$ for $i$-th token at $\tau$-th transformer block.
@@
4 Experiments
4.1 Datasets
We evaluate HyperMoE on 20 NLP datasets across diverse tasks, including sequence classification, question answering, summarization, and text generation. GLUE (Wang et al.,, 2018) and SuperGLUE (Wang et al.,, 2019) benchmarks are widely used evaluation datasets for assessing natural language understanding capabilities. Both of them are a collection of text classification tasks: sentence similarity (STS-B; Cer et al.,, 2017), (MRPC; Dolan and Brockett,, 2005), (QQP; Wang et al.,, 2018), question-answering (BoolQ; Clark et al.,, 2019), (MultiRC; Khashabi et al.,, 2018), (RECORD; Zhang et al.,, 2018), sentiment analysis (SST-2; Socher et al.,, 2013), sentence acceptability (CoLA; Warstadt et al.,, 2019), natural language inference (MNLI; Williams et al.,, 2018), (QNLI; Demszky et al.,, 2018), (RTE; Giampiccolo et al.,, 2007), (CB; De Marneffe et al.,, 2019), word sense disambiguation (WIC; Pilehvar and Camacho-Collados,, 2019), coreference resolution (WSC; Levesque et al.,, 2012) and sentence completion (COPA; Roemmele et al.,, 2011). For the question-answering task, we consider SQuAD v1.1 (Rajpurkar et al.,, 2016), a collection of question-answer pairs derived from Wikipedia articles, with each answer being a text span from the corresponding reading passage. For the summarization task, we use Xsum (Narayan et al.,, 2018) and CNN/Daily Mail(CNNDM) (Hermann et al.,, 2015) to test the model’s ability to summarize articles. And finally, the WikiText-2 dataset (Merity et al.,, 2017) is used to measure the ability of long-range dependencies generation.
4.2 Experiments Details
Following (He et al.,, 2023), we fine-tune pre-trained MoE models on downstream tasks and report results from the last checkpoint. Unless otherwise specified, Our base model primarily uses Switch Transformer-base-8, which is an MoE model built on T5-base (Raffel et al.,, 2020) with 8 available experts, having a total number of parameters of 620M. For the WikiText dataset, we employ GPT-2 small (Radford et al.,, 2019) as the base model and expand it into the MoE structure by duplicating the weights of the feed-forward layer. In addition, we also use Switch Transformer-base-16/32 to explore the effect of expert numbers on our method. To achieve a fair comparison, all methods in our paper employ the same Top-1 routing and auxiliary loss. For different data scales, we grid-search the training epoch and batch size from {10, 15, 20}, and {8, 16, 32, 64}, respectively. The learning rate is grid-search from {1e-5, 5e-5, 1e-4, 5e-4} with Adam optimizer and the first 10% warm-up steps. We set the maximum token length to 1024 for WikiText datasets, 348 for SQuAD, and 256 for all other datasets except for the summarization task. For Xsum and CNNDM, we set the max length of source articles to be 1024 and the max length of the target summary to be 128. As for All experiments run for 3 times with different seeds, and we report the average for each result.
4.3 Baselines
Our approach is built upon Switch Transformer (Fedus et al.,, 2022), a well-known MoE model using Top-1 routing. Consequently, we primarily compare our approach with the following baselines: (1) MoE, fully finetuning switch transformer model. (2) MoE-Share, as it is a relevant baseline that does not exploit the inductive bias of the relationship between selected and unselected experts in the process of computation: add an MLP network that is shared among all experts in the MoE layer of a switch transformer, which has the same size as the experts in MoE.
4.4 Results and Analysis
4.4.1 Main Results
@@ 4.4.2 Ablation Study
We conduct an ablation study on the SQuAD to evaluate the effectiveness of the proposed modules. The embedding design is removed to verify the effect of using external information as embeddings. As shown in Table 3 (row 1), when the embedding and hypernet are removed, our method is equivalent to MoE. Table 3 (row 2) omits the embedding design, directly using the sample’s hidden state as input to the hypernetwork. This results in a marked decrease in performance, even falling below that of MoE. This suggests that conditioning the hypernetwork on the sample enlarges the parameter search space and is difficult to optimize. In an additional experiment, we use a depthwise separable convolutional network (Howard et al.,, 2017) with kernels of sizes 5×5 and 3×3 to compress and reduce the dimensions of the experts’ weights, obtaining expert embeddings. More details are in Appendix A. The selection embeddings are then computed and input into the hypernetwork as described in Section 3.2. Empirically, expert weights can better represent the information of experts. However, as shown in Table 3 (row 3), this strategy leads to a slight drop in performance, defying expectations. A potential explanation is the substantial information loss associated with compressing expert weights, resulting in a loss of specific information details. We leave the exploration of this strategy to future work.
@@
4.4.3 Performance in Scaling the Number of Experts
To explore the impact of the variation in the number of experts on our method, we fine-tuned on the SQuAD dataset using Switch Transformer-base-16/32 as pre-trained models. These models possess 16 and 32 experts in each MoE layer, respectively. As demonstrated in Figure  3, every method achieves performance enhancement across models featuring a diverse number of experts. Notably, our method exhibits consistent superior growth and outperforms the others. This indicates that the proposed conditional generation strategy can still effectively benefit from knowledge transfer as the number of experts increases.
@@ 4.4.3 Performance in Scaling the Number of Experts
Expert embeddings and the selection embeddings have a corresponding relationship. In addition, to explore whether the embeddings encode the information in our proposed method, we provide visualizations of the expert embeddings and computed selection embeddings within the final MoE layer of Switch Transformer-base-8 learned on CNNDM. Figure 5 reveals that both sets of embeddings exhibit sparse distributions, suggesting that the embeddings encode some specific non-relevant information. We also observe a correlation between the distances among selection embeddings and the distances among expert embeddings, such as between 4-5-6, 1-2-8, and 7-3. This correlation implies that the information of the unselected experts encoded by the selection embeddings depends on the information of the selected experts, further illustrating that the selection embeddings effectively capture the information of the knowledge the currently selected experts need.

### v3

@@ Abstract
The Mixture of Experts (MoE) for language models has been proven effective in augmenting the capacity of models by dynamically routing each input token to a specific subset of experts for processing. Despite the success, most existing methods face a challenge for balance between sparsity and the availability of expert knowledge: enhancing performance through increased use of expert knowledge often results in diminishing sparsity during expert selection. To mitigate this contradiction, we propose HyperMoE, a novel MoE framework built upon Hypernetworks. This framework integrates the computational processes of MoE with the concept of knowledge transferring in multi-task learning. Specific modules generated based on the information of unselected experts serve as supplementary information, which allows the knowledge of experts not selected to be used while maintaining selection sparsity. Our comprehensive empirical evaluations across multiple datasets and backbones establish that HyperMoE significantly outperforms existing MoE methods under identical conditions concerning the number of experts. Our code is publicly available at https://github.com/Bumble666/Hyper_MoE
HyperMoE: Towards Better Mixture of Experts via Transferring Among Experts
Hao Zhao1 Zihan Qiu3 Huijia Wu1 Zili Wang4 Zhaofeng He1{NoHyper}thanks:   Corresponding authors. Jie Fu211footnotemark: 1 1Beijing University of Posts and Telecommunications 2Hong Kong University of Science and Technology 3Tsinghua University  4INF Technology {haozhao,huijiawu,zhaofenghe}@bupt.edu.cnjiefu@ust.hk qzh11628@gmail.comziliwang.do.gmail.com
@@ 1 Introduction
The accelerated advancement of large language models has culminated in their widespread application across various domains, including healthcare, education, and social interactions (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023). The remarkable capabilities of these models are attributed to the enhancements in their scale. Nevertheless, the scaling of dense models is often hampered by significant computational demands, posing a challenge to developing the Natural Language Processing (NLP) community. In response, sparse activation models have emerged as a solution (Artetxe et al., 2022; Du et al., 2022), activating only a subset of parameters for different inputs, thus mitigating computational costs. One of the most representative methods is the Mixture of Experts (MoE, Shazeer et al. (2017)), which routers different inputs to specific groups of experts, thereby enlarging the model’s capacity without increasing computational burdens.
@@
The key to effectively reducing computational costs lies in the sparsity of expert selection, with the number of experts selected for each token being kept at a lower level. In practical applications or experiments, existing works (Roller et al., 2021a; Fedus et al., 2022; Rajbhandari et al., 2022; Xue et al., 2023) usually select only one or two experts per input. However, increasing the number of selected experts per token can enhance the availability of expert knowledge and improve the performance of downstream tasks Yang et al. (2019); Shazeer et al. (2017); He et al. (2023). This scenario positions MoE model in a predicament akin to a zero-sum game: a choice between increasing the number of available experts to improve performance or preserving a lower level of available experts to ensure sparsity, as depicted in Figure 1.
To mitigate this contradiction, one solution would be to use the knowledge of other experts to assist the sparsely selected experts. This is similar to multi-task learning, which transfers knowledge among related tasks. Some works (Karimi Mahabadi et al., 2021; Ivison and Peters, 2022; Zhao et al., 2023) suggest using hypernetworks (Ha et al., 2017) to generate task-specific knowledge to enhance positive transfer between tasks. Inspired by this, we aim to increase the availability of expert knowledge by transferring the knowledge of unselected experts while sparsely selecting experts.
@@
We further improve upon this by introducing the concept of cross-layer Hypernetworks: A hypernetwork is shared among all transformer layers, which enables information flow among MoEs in different layers. This brings additional efficiency in terms of parameters and computational costs: Despite the additional computation, our method only experienced a decrease [^1] The degree of decline in speed is related to the scale of the Hypernetworks and the bottleneck size in the generated HyperExpert (similar to $r$ in LoRA). For various tasks, these hyperparameters can be dynamically adjusted to control the delay. of approximately 15% in training speed and 10% in inference speed compared to the standard MoE.
We evaluate HyperMoE on 20 representative NLP datasets across diverse tasks: sequence classification, extractive question answering, summarization, and text generation. Extensive experimental results show that HyperMoE outperforms baselines, including Switch Transformer (Fedus et al., 2022) with MoE architecture. This demonstrates the effectiveness of our method in transferring knowledge to experts, which increases the utilization of expert knowledge while keeping the number of experts selected at a low level.
@@
  • We propose a novel HyperMoE architecture with HyperExpert for MoE framework, which resolves the inherent tension between maintaining sparse expert selection and ensuring sufficient expert availability within MoE.
  • HyperMoE outperforms baselines based on Switch Transformer across a diverse set of NLP tasks, confirming our approach’s effectiveness.
@@ 2.1 Mixture of Expert
The gate model $G(\cdot)$ is a Noisy Top-K Network Shazeer et al. (2017) with parameters $W_{g}$ and $W_{noise}$. This gating method introduces adjustable noise and then retains the top-k values as the final output:
@@ 2.2 HyperNetworks
Hypernetwork (Ha et al., 2017) can generate parameters to be used by target networks or modules. Specifically, a hypernetwork with independent parameters $\phi$ denoted as $h_{\phi}$, leverages an context information $z$ to generate the target parameters $\theta$ for the primary network $f_{\theta}$ and the primary network with an input $x$ is redefined as:
@@
This method of flexibly adjusting the parameters of the target network to adapt to different input scenarios is widely used in multi-task learning (Karimi Mahabadi et al., 2021; Üstün et al., 2022) and few-shot learning (Ponti et al., 2021). While generating condition-specific parameters, these parameters also benefit from shared knowledge (Pfeiffer et al., 2023).
@@ 3.1 Conditional Expert
In our approach, the matrices $W_{1}$ and $W_{2}$ are generated by a hypernetwork as described in Section 3.2. In addition, we adopt a bottleneck structure for the conditional expert to improve parameter efficiency inspired by the Adapter (Houlsby et al., 2019). Specifically, the bottleneck dimension $b$ satisfies $b\ll h$ in our method.
@@ 3.2 HyperExpert
These works (Karimi Mahabadi et al., 2021; He et al., 2022; Phang et al., 2023; Ivison et al., 2023) indicate that hypernetworks can learn the parameter information of the main neural network under different input scenarios and efficiently adjust the parameters of the target network to adapt to this information.
@@
The weight matrices of hypernetworks ${W^{D}}$ and ${W^{U}}$ are used to generate the down-projection matrix ${D}_{i}^{\tau}\in\mathbb{R}^{h\times b}$ and the up-projection matrix ${U}_{i}^{\tau}\in\mathbb{R}^{b\times h}$ in the HyperExpert $\hat{E}_{i}$ for $i$-th token at $\tau$-th transformer block.
@@
4 Experiments
4.1 Datasets
We evaluate HyperMoE on 20 NLP datasets across diverse tasks including sequence classification, question answering, summarization, and text generation. GLUE (Wang et al., 2018) and SuperGLUE (Wang et al., 2019) benchmarks are widely used evaluation datasets for assessing natural language understanding capabilities. Both of them are a collection of text classification tasks: sentence similarity (STS-B; Cer et al., 2017), (MRPC; Dolan and Brockett, 2005), (QQP; Wang et al., 2018), question-answering (BoolQ; Clark et al., 2019), (MultiRC; Khashabi et al., 2018), (RECORD; Zhang et al., 2018), sentiment analysis (SST-2; Socher et al., 2013), sentence acceptability (CoLA; Warstadt et al., 2019), natural language inference (MNLI; Williams et al., 2018), (QNLI; Demszky et al., 2018), (RTE; Giampiccolo et al., 2007), (CB; De Marneffe et al., 2019), word sense disambiguation (WIC; Pilehvar and Camacho-Collados, 2019), coreference resolution (WSC; Levesque et al., 2012) and sentence completion (COPA; Roemmele et al., 2011). For the question-answering task, we consider SQuAD v1.1 (Rajpurkar et al., 2016), a collection of question-answer pairs derived from Wikipedia articles, with each answer being a text span from the corresponding reading passage. For the summarization task, we use Xsum (Narayan et al., 2018) and CNN/Daily Mail(CNNDM) (Hermann et al., 2015) to test the model’s ability to summarize articles. And finally, the WikiText-2 dataset (Merity et al., 2017) is used to measure the ability of long-range dependencies generation.
4.2 Experiments Details
Following (He et al., 2023), we fine-tune pre-trained MoE models on downstream tasks and report results from the last checkpoint. Unless otherwise specified, Our base model primarily uses Switch Transformer-base-8, which is an MoE model built on T5-base (Raffel et al., 2020) with 8 available experts, having a total number of parameters of 620M. For the WikiText dataset, we employ GPT-2 small (Radford et al., 2019) as the base model and expand it into the MoE structure by duplicating the weights of the feed-forward layer. In addition, we also use Switch Transformer-base-16/32 to explore the effect of expert numbers on our method. To achieve a fair comparison, all methods in our paper employ the same Top-1 routing and auxiliary loss. For different data scales, we grid-search the training epoch and batch size from {10, 15, 20}, and {8, 16, 32, 64}, respectively. The learning rate is grid-search from {1e-5, 5e-5, 1e-4, 5e-4} with Adam optimizer and the first 10% warm-up steps. We set the maximum token length to 1024 for WikiText datasets, 348 for SQuAD, and 256 for all other datasets except for the summarization task. For Xsum and CNNDM, we set the max length of source articles to be 1024 and the max length of the target summary to be 128. As for All experiments run for 3 times with different seeds and we report the average for each result.
4.3 Baselines
Our approach is built upon Switch Transformer (Fedus et al., 2022), a well-known MoE model using Top-1 routing. Consequently, we primarily compare our approach with the following baselines: (1) MoE, fully finetuning switch transformer model. (2) MoE-Share, as it is a relevant baseline that does not exploit the inductive bias of the relationship between selected and unselected experts in the process of computation: add an MLP network that is shared among all experts in the MoE layer of a switch transformer, which has the same size as the experts in MoE.
4.4 Results and Analysis
4.4.1 Main Results
@@
Additionally, we compared our method with a related baseline (Do et al., 2023) that uses a hypernetwork to generate conditional parameters for the routing of MoEs, aiming to address the routing allocation problem in SMoE methods, where all experts tend to have similar representations. Experimental results can be found in Appendix B.
@@ 4.4.2 Ablation Study
We conduct an ablation study on the SQuAD to evaluate the effectiveness of the proposed modules. The embedding design is removed to verify the effect of using external information as embeddings. As shown in Table 3 (row 1), when the embedding and hypernet are removed, our method is equivalent to MoE. Table 3 (row 2) omits the embedding design, directly using the sample’s hidden state as input to the hypernetwork. This results in a marked decrease in performance, even falling below that of MoE. This suggests that conditioning the hypernetwork on the sample enlarges the parameter search space and is difficult to optimize. In an additional experiment, we use a depthwise separable convolutional network (Howard et al., 2017) with kernels of sizes 5×5 and 3×3 to compress and reduce the dimensions of the experts’ weights, obtaining expert embeddings. More details are in Appendix A. The selection embeddings are then computed and input into the hypernetwork as described in Section 3.2. Empirically, expert weights can better represent the information of experts. However, as shown in Table 3 (row 3), this strategy leads to a slight drop in performance, defying expectations. A potential explanation is the substantial information loss associated with compressing expert weights, resulting in a loss of specific information details. We leave the exploration of this strategy to future work.
@@
4.4.3 Performance in Scaling the Number of Experts and Model Size
To explore the impact of the variation in the number of experts on our method, we fine-tuned on the SQuAD dataset using Switch Transformer-base-16/32 as pre-trained models. These models possess 16 and 32 experts in each MoE layer, respectively. As demonstrated in Figure 3, every method achieves performance enhancement across models featuring a diverse number of experts. Notably, our method exhibits consistent superior growth and outperforms the others. This indicates that the proposed conditional generation strategy can still effectively benefit from knowledge transfer as the number of experts increases.
@@ 4.4.3 Performance in Scaling the Number of Experts and Model Size
In addition, to verify the effectiveness of our method on larger parameter-scale language models, we fine-tuned DeepSeek-MoE (Dai et al., 2024), which has 16 billion parameters, on the Open-Platypus (Lee et al., 2023) dataset and tested the fine-tuned model on MMLU (Hendrycks et al., 2020), GSM8K (Cobbe et al., 2021), and BBH (Suzgun et al., 2023). As shown in Table 4, our method consistently achieved improvements. This indicates that our method can be widely applied to language models ranging from hundreds of millions to tens of billions of parameters.
Method MMLU GSM8K BBH
MoE 46.52 17.97 39.38
HyperMoE 47.06 18.75 40.77
Table 4: Comparison of results after fine-tuning on DeepSeek-MoE 16B.
@@
Expert embeddings and the selection embeddings have a corresponding relationship. In addition, to explore whether the embeddings encode the information in our proposed method, we provide visualizations of the expert embeddings and computed selection embeddings within the final MoE layer of Switch Transformer-base-8 learned on CNNDM. Figure 5 reveals that both sets of embeddings exhibit sparse distributions, suggesting that the embeddings encode some specific non-relevant information. We also observe a correlation between the distances among selection embeddings and the distances among expert embeddings, such as between 4-5-6, 1-2-8, 7-3. This correlation implies that the information of the unselected experts encoded by the selection embeddings depends on the information of the selected experts, further illustrating that the selection embeddings effectively capture the information of the knowledge the currently selected experts need.